A cathode ray tube (CRT) is used to display images in some televisions and computer monitors by accelerating electrons from the cathode toward the anode, which is at a higher potential than the cathode. If the potential difference between the cathode and anode is 25 kV, at what speed do the electrons move as they reach the anode? Express your answer in scientific notation. (in m/s) (Hint: energy is conserved)
Answer #1

The electrons are accelereated through potential difference of 25kV, hence

work done = energy = qV = 25keV

But energy = kinetic energy of electron = 1/2mv^2

1/2mv^2 = 25*1000*1.6*10^-19

1/2 * 9.1 * 10^-31*v^2 =25*1000*1.6*10^-19

Hence v = 9.4 x 10^7 m/s
Answer #2

The kinetic energy is acquired by the electron from the potential difference:

K=\frac{1}{2}m_e.v^2=q_e.\Delta V (1)

Where

m_e=9.1.10^{-31}Kg

q_e=1.6.10^{-19}C

Solving for "v" in (1):

v=\sqrt{\frac{2q_e.\Delta V}{m_e}}\Rightarrow \mathbf{v=9.4.10^7m/s}